# GD-VAEs: Geometric Dynamic Variational Autoencoders for Learning Nonlinear Dynamics and Dimension Reductions

### v3

@@ 2 Learning with Manifold Latent Spaces having General Geometries and Topologies
From equations 24 and 25, the gradients $\nabla_{\mathbf{u}}G$, $\nabla_{\mathbf{w}}G$, and using equation 22, we compute $\nabla_{\phi}\mathbf{z}^{*}$. This allows us to learn VAEs with latent spaces for $\mathbf{z}$ with general specified topologies and controllable geometric structures. We provide additional details in Appendix A. We refer to this learning framework for data-driven modeling of dynamics on latent spaces having general geometries as Geometric Dynamic Variational Autoencoders (GD-VAEs).
@@ 3.1 Burgers’ Equation of Fluid Mechanics: Learning Nonlinear PDE Dynamics
We train the model on the pairs $(u(x,t),u(x,t+\tau))$ by drawing $m$ samples of $u^{k_{i}}(x,t_{i})\in\mathcal{S}_{t_{i}}\mathcal{U}$ which generates the evolved state under Burgers equation $u^{k_{i}}(x,t_{i}+\tau)$ over time-scale $\tau$. We perform studies with parameters $\nu=2\times 10^{-2}$, $\tau=2.5\times 10^{-1}$ with our Deep Neural Networks (DNNs) with layer sizes (in)-400-400-(out), ReLU activations, and $\gamma=0.5$, $\beta=1$. The covariances $\boldsymbol{\Sigma}_{e}$ and $\boldsymbol{\Sigma}_{d}$ used are diagonal with initial standard deviations $\sigma_{d}=\sigma_{e}=4\times 10^{-3}$ so that $\boldsymbol{\Sigma}_{e}=\mbox{diag}(\sigma_{e}^{2})$ and $\boldsymbol{\Sigma}_{d}=\mbox{diag}(\sigma_{d}^{2})$. We show results of our model predictions in Figure 4 and Table 1.
Figure 4: Burgers’ Equation: Prediction of Dynamics. We consider responses for initial conditions $\mathcal{U}_{1}=\{u\,|\,u(x,t;\alpha)=\alpha\sin(2\pi x)+(1-\alpha)\cos^{3}(2\pi x)\}$. Predictions are made for the evolution $u$ over the time-scale $\tau$ satisfying equation 26 with initial conditions in $\mathcal{U}_{1}$. We find our nonlinear VAE methods are able to learn with $2$ latent dimensions the dynamics with errors $<1\%$. Methods such as DMD [74; 81] with $3$ modes which are only able to use a single linear space to approximate the initial conditions and prediction encounter challenges in approximating the nonlinear evolution. We find our linear VAE method with $2$ modes provides some improvements by allowing for using different linear spaces for representing the input and output functions. Results are summarized in Table 1.
@@
We investigate the importance of the non-linear approximation properties of our methods in capturing system behaviors. We do this by making comparisons with linear methods that include Dynamic Mode Decomposition (DMD) [74; 81], Proper Orthogonal Decomposition (POD) [15], and a linear variant of our GD-VAE approach. Recent CNN-AEs have also studied related advantages of non-linear approximations [51]. Some distinctions in our work is the use of VAEs to further regularize the autoencoders and using topological latent spaces to facilitate further capturing inherent structures. The DMD and POD are widely used and successful approaches that aim to find an optimal linear space on which to project the dynamics and learn a linear evolution law for system behaviors. DMD and POD have been successful in obtaining models for many applications, including steady-state fluid mechanics and transport problems [81; 74]. However, given their inherent linear approximations they can encounter well-known challenges related to translational and rotational invariances, as arise in advective phenomena and other settings [10]. Our comparison studies can be found in Table 1.
@@ 3.3 Periodic and Doubly-Periodic Systems: Cylinder and Torus Latent Spaces
This corresponds to the topology of a torus. We can project our solutions onto the Clifford Torus given by the product space $S^{1}\times S^{1}\subset\mathbb{R}^{4}$, where $S^{1}$ is the circle space [44; 83]. For example representations of latent spaces, see Figure 6. In each case, we seek representations that disentangle state information of the collection of solutions $\{u_{\alpha}\}_{0\leq\alpha\leq 1}$ or $\{u_{\alpha_{1},\alpha_{2}}\}_{0\leq\alpha_{1},\alpha_{2}\leq 2\pi}$.
@@
This mapping corresponds in a latent space having a cylinder topology. Similarly, for the doubly-periodic family $\{u_{\alpha_{1},\alpha_{2}}\}$ the state information can be mapped in principle to a torus by $\mathbf{f}(\boldsymbol{\alpha})$. This would yield the disentangled embedding $u_{\boldsymbol{\alpha}}(x,t)\rightarrow\left(\mathbf{f}(\boldsymbol{\alpha}),t\right)$ with $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2})$. The torus can be represented in principle in $\mathbb{R}^{3}$ or as a product space $S^{1}\times S^{1}$ where $S^{1}$ is the circle space. This latter representation is referred to as the Clifford Torus [44; 83]. This gives the latent space with the cylinder topology $\mathcal{B}\times\mathbb{R}$, where $\mathcal{B}$ is the base-space. For periodic case the $\mathcal{B}=S^{1}$ is a circle and for the doubly-periodic case the $\mathcal{B}=S^{1}\times S^{1}$ is a torus.
@@ 3.4 Constrained Mechanical Systems: Learning with General Manifold Latent Spaces
We consider physical systems with constrained mechanics, such as the arm mechanism for reaching for objects in Figure 8. The observations are taken to be the two locations $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathbb{R}^{2}$ giving $\mathbf{x}=(\mathbf{x}_{1},\mathbf{x}_{2})\in\mathbb{R}^{4}$. When the segments are constrained, these configurations lie on a manifold embedded in $\mathbb{R}^{4}$. The aim of these studies is to show how GD-VAEs with manifold latent spaces can be used to learn representations for constrained systems. We do not currently consider dynamical predictions in these studies. This corresponds to GD-VAEs with latent space dynamics $\mathbf{z}^{\prime}=f_{\theta_{\ell}}(\mathbf{z})=\mathbf{z}$ given by the identity map. We focus instead on the types of representations learned for the constrained latent geometries. When the two segments are constrained to be rigid, this results in the collection of configurations having the topology of a torus. We can also allow the segments to extend and consider more exotic constraints. For example, we could require the two points $\mathbf{x}_{1},\mathbf{x}_{2}$ always be within a constraint set that is a klein bottle surface within $\mathbb{R}^{4}$. Related situations arise in other areas of imaging and mechanics, such as in pose estimation and in studies of visual perception  [64; 13; 72].
@@ 3.5 Reaction-Diffusion PDEs in 2D: Learning Representations for Spatially Distributed Dynamics
The $u=u(x,t)$ and $v(x,t)$ give the spatially distributed concentration of each chemical species at time $t$ with $x\in\mathbb{R}^{2}$. We consider the case with periodic boundary conditions with $x\in[0,L]\times[0,L]$. We develop learning methods for investigating the Brusselator system  [67; 66], which is known to have regimes exhibiting limit cycles [77; 35; 36]. This indicates after an initial transient, the orbit of the dynamics will localize and approach a subset of states topologically similar to a circle. We show how GD-VAE can utilize this topological information to construct latent spaces for encoding states of the system. The Brusselator [67; 66] has reactions with $f(u,v)=a-(1+b)u+vu^{2}$ and $g(u,v)=bu-vu^{2}$. We take throughout the diffusivity $D_{1}=1,D_{2}=0.1$ and reaction rates $a=1,b=3$.
@@
We have developed software packages for the introduced GD-VAE methods. This includes python implementations of our geometric projections (g-projection), manifold latent space representations, custom back-propagation approaches for modular training, and examples. The python package can be installed using pip install gd-vae-pytorch or downloaded from our website. For more information, see http://atzberger.org/.
@@ Acknowledgements
Authors research supported by grants DOE Grant ASCR PHILMS DE-SC0019246 and NSF Grant DMS-1616353. Also to R.N.L. support by a donor to UCSB CCS SURF program. Authors also acknowledge UCSB Center for Scientific Computing NSF MRSEC (DMR1121053) and UCSB MRL NSF CNS-1725797. P.J.A. would also like to acknowledge a hardware grant from Nvidia.
@@ Appendix C Identifying Latent Geometric Structures using VAE Covariance $\boldsymbol{\Sigma}_{e}$
Figure 15: Extracting Latent Geometric Structure using Learned VAE Variances. We use the $\sigma_{e,i}$ and $\mu_{e,i}$ of VAEs to further extract geometric information from data and to perform reductions similar to [54]. By looking at mini-batches of data, we can compute two statistics (i) the mean of the variance (mov) $q_{i,mov}$ and (ii) the variance of the mean (vom) $q_{i,vom}$, see equations 36 and 37. We show results for Burgers’ Equation with embedding in an unconstrained $10$ dimensional space. We see the most relevant feature dimensions are the ones with $q_{i,vom}$ large and $q_{i,mov}$ small (left). This yields a reduced three dimensional embedding with a good representation for the state information using $z_{1},z_{9}$ coordinates (middle). In contrast, using two arbitary dimensions such as coordinates $z_{1},z_{2}$ yields a poor representation which is entangled and mapped to small scales (right). This provides further ways to extract geometric information from the data and representations of latent structures.

### v4

@@ 2 Learning with Manifold Latent Spaces having General Geometries and Topologies
From equations 25 and 26, the gradients $\nabla_{\mathbf{u}}G$, $\nabla_{\mathbf{w}}G$, and using equation 23, we compute $\nabla_{\phi}\mathbf{z}^{*}$. This allows us to learn VAEs with latent spaces for $\mathbf{z}$ with general specified topologies and controllable geometric structures. We provide additional details in Appendix A. We refer to this learning framework for data-driven modeling of dynamics on latent spaces having general geometries as Geometric Dynamic Variational Autoencoders (GD-VAEs).
@@ 3.1 Burgers’ Equation of Fluid Mechanics: Learning Nonlinear PDE Dynamics
We train the model by drawing $m$ samples of the prediction pairs $(u^{k_{i}}(x,t_{i}),u^{k_{i}}(x,t_{i}+\tau))$. These are sampled from solution trajectories with $u^{k}(x,t)\in\mathcal{S}_{t_{i}}\mathcal{U}$. The $\mathcal{U}$ gives the ensemble of initial conditions. The $\mathcal{S}_{t}$ denotes the non-linear solution operator for the Burgers equation with $u(x,t)=\mathcal{S}_{t}[u(x,0)]$. For the $i^{th}$ sample, the $k_{i}$ gives the trajectory index $k$ and the $t_{i}$ the sampled initial time $t$ from the trajectory. We perform studies with parameters $\nu=2\times 10^{-2}$, $\tau=2.5\times 10^{-1},m=100$ with our Deep Neural Networks (DNNs) with layer sizes (in)-400-400-(out), ReLU activations, and $\gamma=0.5$, $\beta=1$. The covariances $\boldsymbol{\Sigma}_{e}$ and $\boldsymbol{\Sigma}_{d}$ used are diagonal with initial standard deviations $\sigma_{d}=\sigma_{e}=4\times 10^{-3}$ so that $\boldsymbol{\Sigma}_{e}=\mbox{diag}(\sigma_{e}^{2})$ and $\boldsymbol{\Sigma}_{d}=\mbox{diag}(\sigma_{d}^{2})$. We show results of our model predictions in Figure 4 and Table 1.
Figure 4: Burgers’ Equation: Prediction of Dynamics. We consider responses for initial conditions $\mathcal{U}_{1}=\{u\,|\,u(x,t;\alpha)=\alpha\sin(2\pi x)+(1-\alpha)\cos^{3}(2\pi x)\}$. Predictions are made for the evolution $u$ over the time-scale $\tau$ satisfying equation 27 with initial conditions in $\mathcal{U}_{1}$. We find our nonlinear VAE methods are able to learn with $2$ latent dimensions the dynamics with errors $<1\%$. Methods such as DMD [76; 84] with $3$ modes which are only able to use a single linear space to approximate the initial conditions and prediction encounter challenges in approximating the nonlinear evolution. We find our linear VAE method with $2$ modes provides some improvements by allowing for using different linear spaces for representing the input and output functions. Results are summarized in Table 1.
@@
We investigate the importance of the non-linear approximation properties of our methods in capturing system behaviors. We do this by making comparisons with linear methods that include Dynamic Mode Decomposition (DMD) [76; 84], Proper Orthogonal Decomposition (POD) [15], and a linear variant of our GD-VAE approach. Recent CNN-AEs have also studied related advantages of non-linear approximations [51]. Some distinctions in our work is the use of VAEs to further regularize the autoencoders and using topological latent spaces to facilitate further capturing inherent structures. The DMD and POD are widely used and successful approaches that aim to find an optimal linear space on which to project the dynamics and learn a linear evolution law for system behaviors. DMD and POD have been successful in obtaining models for many applications, including steady-state fluid mechanics and transport problems [84; 76]. However, given their inherent linear approximations they can encounter well-known challenges related to translational and rotational invariances, as arise in advective phenomena and other settings [10]. Our comparison studies can be found in Table 1.
@@ 3.3 Periodic and Doubly-Periodic Systems: Cylinder and Torus Latent Spaces
This corresponds to the topology of a torus. We can project our solutions onto the Clifford Torus given by the product space $S^{1}\times S^{1}\subset\mathbb{R}^{4}$, where $S^{1}$ is the circle space [44; 86]. For example representations of latent spaces, see Figure 6. In each case, we seek representations that disentangle state information of the collection of solutions $\{u_{\alpha}\}_{0\leq\alpha\leq 1}$ or $\{u_{\alpha_{1},\alpha_{2}}\}_{0\leq\alpha_{1},\alpha_{2}\leq 2\pi}$.
@@
This mapping corresponds in a latent space having a cylinder topology. Similarly, for the doubly-periodic family $\{u_{\alpha_{1},\alpha_{2}}\}$ the state information can be mapped in principle to a torus by $\mathbf{f}(\boldsymbol{\alpha})$. This would yield the disentangled embedding $u_{\boldsymbol{\alpha}}(x,t)\rightarrow\left(\mathbf{f}(\boldsymbol{\alpha}),t\right)$ with $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2})$. The torus can be represented in principle in $\mathbb{R}^{3}$ or as a product space $S^{1}\times S^{1}$ where $S^{1}$ is the circle space. This latter representation is referred to as the Clifford Torus [44; 86]. This gives the latent space with the cylinder topology $\mathcal{B}\times\mathbb{R}$, where $\mathcal{B}$ is the base-space. For periodic case the $\mathcal{B}=S^{1}$ is a circle and for the doubly-periodic case the $\mathcal{B}=S^{1}\times S^{1}$ is a torus.
@@ 3.4 Constrained Mechanical Systems: Learning with General Manifold Latent Spaces
We consider physical systems with constrained mechanics, such as the arm mechanism for reaching for objects in Figure 8. The observations are taken to be the two locations $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathbb{R}^{2}$ giving $\mathbf{x}=(\mathbf{x}_{1},\mathbf{x}_{2})\in\mathbb{R}^{4}$. When the segments are constrained, these configurations lie on a manifold embedded in $\mathbb{R}^{4}$. The aim of these studies is to show how GD-VAEs with manifold latent spaces can be used to learn representations for constrained systems. We do not currently consider dynamical predictions in these studies. This corresponds to GD-VAEs with latent space dynamics $\mathbf{z}^{\prime}=f_{\theta_{\ell}}(\mathbf{z})=\mathbf{z}$ given by the identity map. We focus instead on the types of representations learned for the constrained latent geometries. When the two segments are constrained to be rigid, this results in the collection of configurations having the topology of a torus. We can also allow the segments to extend and consider more exotic constraints. For example, we could require the two points $\mathbf{x}_{1},\mathbf{x}_{2}$ always be within a constraint set that is a klein bottle surface within $\mathbb{R}^{4}$. Related situations arise in other areas of imaging and mechanics, such as in pose estimation and in studies of visual perception  [64; 13; 74].
@@ 3.5 Reaction-Diffusion PDEs in 2D: Learning Representations for Spatially Distributed Dynamics
The $u=u(x,t)$ and $v(x,t)$ give the spatially distributed concentration of each chemical species at time $t$ with $x\in\mathbb{R}^{2}$. We consider the case with periodic boundary conditions with $x\in[0,L]\times[0,L]$. We develop learning methods for investigating the Brusselator system  [67; 66], which is known to have regimes exhibiting limit cycles [80; 35; 36]. This indicates after an initial transient, the orbit of the dynamics will localize and approach a subset of states topologically similar to a circle. We show how GD-VAE can utilize this topological information to construct latent spaces for encoding states of the system. The Brusselator [67; 66] has reactions with $f(u,v)=a-(1+b)u+vu^{2}$ and $g(u,v)=bu-vu^{2}$. We take throughout the diffusivity $D_{1}=1,D_{2}=0.1$ and reaction rates $a=1,b=3$.
@@
We have developed software packages for the introduced GD-VAE methods. This includes python implementations of our geometric projections (g-projection), manifold latent space representations, custom back-propagation approaches for modular training, and examples. The python package can be installed using pip install gd-vae-pytorch or downloaded from our website. For more information, see https://atzberger.org/.
@@ Acknowledgements
Authors research supported by grants DOE Grant ASCR PHILMS DE-SC0019246, NSF Grant DMS-1616353, and NSF-DMS-2306101. Also to R.N.L. support by a donor to UCSB CCS SURF program. Authors also acknowledge UCSB Center for Scientific Computing NSF MRSEC (DMR1121053) and UCSB MRL NSF CNS-1725797. P.J.A. would also like to acknowledge a hardware grant from Nvidia.
@@ Appendix C Identifying Latent Geometric Structures using VAE Covariance $\boldsymbol{\Sigma}_{e}$
Figure 15: Extracting Latent Geometric Structure using Learned VAE Variances. We use the $\sigma_{e,i}$ and $\mu_{e,i}$ of VAEs to further extract geometric information from data and to perform reductions similar to [54]. By looking at mini-batches of data, we can compute two statistics (i) the mean of the variance (mov) $q_{i,mov}$ and (ii) the variance of the mean (vom) $q_{i,vom}$, see equations 37 and 38. We show results for Burgers’ Equation with embedding in an unconstrained $10$ dimensional space. We see the most relevant feature dimensions are the ones with $q_{i,vom}$ large and $q_{i,mov}$ small (left). This yields a reduced three dimensional embedding with a good representation for the state information using $z_{1},z_{9}$ coordinates (middle). In contrast, using two arbitary dimensions such as coordinates $z_{1},z_{2}$ yields a poor representation which is entangled and mapped to small scales (right). This provides further ways to extract geometric information from the data and representations of latent structures.